OpenStudy (anonymous):

prove d/dx(arccosh x) = 1/sqrt(x^2-1) implicitly thanks (:

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[y=\cos^{-1} x\]??

OpenStudy (amistre64):

hyperbolic eh...shudder

OpenStudy (anonymous):

Let arcosh(x) = y cosh(y) = x (d/dx)(cosh(y)) = (d/dx)(x) sinh(y) * (dy/dx) = 1 (dy/dx) = 1/sinh(y) ;;; cosh²(y) - sinh²(y) = 1, ;;; sinh(y) = sqrt(cosh²(y) - 1) = sqrt(x² - 1) Hence, if y = arcosh(x), (d/dx)(arcosh(x)) = dy/dx = 1/sqrt(x² - 1)

OpenStudy (anonymous):

thanks!